if 2 terms in AP series are 4 and 9 in ratio 1:3 .. then find D and A??
Answer by richwmiller(17219)

If you mean that the 4th and 9th terms are in the ratio 1:3
an = a + (n - 1)d
a4 = a + 3d
a9 = a + 8d
(a + 3d)/(a + 8d)=1/3
d = -2a
a=-d/2
a4 = a + 3(-2a)=-5a
a9 = a + 8(-2a)=-15a
let a1=2
d=-4
2,-2,-6,-10,-14,-18
-6/-18=1/3
one answer is
a1=2
d=-4